OpenStudy (anonymous):

write the solution in the set builder notation -3x-8<=-23

OpenStudy (tamtoan):

add -3x to both side , after than add -23 to both side...now what do you get ?

OpenStudy (anonymous):

-11<=-31

OpenStudy (tamtoan):

where's the x ? ..each time you add something to 1 side, you have to add the same thing to the other side of the equation too, can you try that again?

OpenStudy (tamtoan):

sorry, it should be add 3x and 23 ...forget the - sign

OpenStudy (anonymous):

31x

OpenStudy (tamtoan):

ok try this , your inequality is : -3x - 8 <= -23 can you add 3x to both side of the inequality and tell me what you get?

OpenStudy (anonymous):

im so confused

OpenStudy (tamtoan):

add 3x to both side means... your original one is : -3x - 8 <= -23 add 3x to both side is: -3x - 8 +3x <= -23 + 3x now can you simplify this ? what is -3x + 3x = ?

OpenStudy (anonymous):

x

OpenStudy (tamtoan):

what is -3football + 3football ? or -3cups + 3cups? ...x is a varible, think of it like an object.

OpenStudy (anonymous):

0

OpenStudy (tamtoan):

yea, then -3x + 3x = 0 .. so let's write again : -3x - 8 <= -23 -3x - 8 + 3x <= -23 + 3x -8 <= -23 + 3x (since the -3x + 3x = 0 ) now add + 23 to both side of the inequality, can you show me ?

OpenStudy (anonymous):

31

OpenStudy (tamtoan):

it's both side..remember i show you earlier that you add 3x to both side ? to add 23 to both side you have to write like this .... -8 <= -23 + 3x -8 + 23 <- -23 + 3x + 23 can you simplify that ?